Scenario 15-10 Vincent operates a scenic tour business in Boston. He has one bus which can fit 50 people per tour and each tour lasts 2 hours. His total cost of operating one tour is fixed at $450. Vincents cost is not reduced if he runs a tour with a partially full bus. While his cost is the same for all tours, Vincent charges each passenger his/her willingness to pay: adults $18 per trip, children S10 per trip, and senior citizens $12 per trip. At those rates, on a typical day Vincents demand is: Passenger TypeWIllingness to Pay Demand per day 18 2s $12 Assume that Vincents customers are always available for the tour, therefore, he can fill his bus for each tour as long as there is sufficient total demand for the day. Refer to Scenario 15-10. What is Vincents total revenue on a typical day? Oa. $1,500 b. S1,800 с. S2,170 d.$2,700
Answer #1

Answer
Total demand =70+25+55=150
Vincent can take 50 at a time so he will make three trips of 2 hours for that day
the total revenue =price *quantity
=70*18+10*25+12*55
=2170
the total revenue for a typical day is $2170
option c